What is the molarity of a solution containing 33.5 g of NaCl dissolved in 500. mL of water?


1
Expert's answer
2022-05-16T17:10:04-0400

Solution:

The molar mass of NaCl is 58.44 g/mol

Therefore,

Moles of NaCl = (33.5 g NaCl) × (1 mol NaCl / 58.44 g NaCl) = 0.573 mol NaCl


The density of water is around approximately 1 g/mL

Therefore,

Mass of water = (500 mL) × (1 g / 1 mL) = 500 g


Mass of solution = Mass of water + Mass of NaCl = 500 g + 33.5 g = 533.5 g

Assume that the density of a solution is 1 g/mL

Therefore,

Liters of solution = (533.5 g) × (1 mL / 1 g) × (1 L / 1000 mL) = 0.5335 L


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of NaCl solution = (0.573 mol) / (0.5335 L) = 1.074 mol/L

Molarity of NaCl solution = 1.074 M


Answer: The molarity of NaCl solution is 1.074 M